OpenStudy (bananas1234):

Which expression makes the equation true? X^8 = (blank)^2 x^4 x^5 x^10 x^12

OpenStudy (anonymous):

do u have any idea?

OpenStudy (bananas1234):

no, sorry, i don't :/

OpenStudy (anonymous):

ok look: (x^n)^2 = x^2n

OpenStudy (anonymous):

\[(x^{a})^{b} = x^{ab} \]

OpenStudy (anonymous):

a * b

OpenStudy (anonymous):

do u understand???

OpenStudy (bananas1234):

kinda, i am still a bit confused

OpenStudy (anonymous):

x^8= (x^???)^2

OpenStudy (anonymous):

u must 8/2

OpenStudy (bananas1234):

ok, so ^4?

OpenStudy (bananas1234):

x^4?

OpenStudy (anonymous):

yes

OpenStudy (bananas1234):

Ok, great!!

OpenStudy (bananas1234):

Do you agree @rational ?

OpenStudy (rational):

looks perfect to me

OpenStudy (bananas1234):

Thanks guys!!